A hydrocarbon has 90%carbon and 10%hydrogen what is its empirical formula

Grade:9

1 Answers

rengapadmanathan
8 years ago
to wfind the simplest atomicity(emperical formula )  of a compound  : first divide the % with its atomic no. you will get a decimal no. now take the least value among the two decimal and divide the decimal with the least decimal value   
90% of carbon => 90/12 =7.5........................ (1)
10% of hydrogen => 10/1 =10 ….....................(2)
here least decimal is 7.5 divide (1) and (2) by 7.5 
 
we get 1 & 4/3 
make it as whole no. so u get 3 & 4 
so the emperical formula is : C3H4